# Narrative Scaffolding: Transforming Data-Driven Sensemaking Through Narrative-First Exploration

### v1

@@ 8.1. The Narrative-First Paradigm: Supporting Natural Reasoning Through Writing
Our findings reveal how narrative construction can function as a primary reasoning interface during data analysis, aligning with narrative psychology’s view that storytelling structures how people make sense of complex information (Murray, 2003; Bruner, 1991). Rather than treating writing as post-analysis documentation (Zhao et al., 2021; Lee et al., 2015), participants used NS not merely to document completed insights, but as a thinking medium, a space where vague hunches evolved into precise, evidence-backed claims through the act of writing.
A key aspect of this narrative-as-thinking-medium approach was adapted to how participants actually think. This paradigm accommodated diverse analytical entry points and supported continuous interpretive refinement throughout exploration. Unlike visualization-first tools that require users to translate thoughts into structured queries (Setlur et al., 2016; Gao et al., 2015), NS accommodated diverse entry points—whether specific hypotheses or exploratory questions. This addresses the ”gulf of envisioning” (Subramonyam et al., 2024): the difficulty of translating vague intentions into system-actionable inputs. Participants could write underspecified hunches, generate visualizations, then iteratively refine their narrative to pursue new directions—each refinement triggering updated views without breaking flow. This fluidity reduced friction not just during early-stage exploration (Wongsuphasawat et al., 2019; Kandel et al., 2012), but throughout the entire sensemaking process, as participants’ understanding evolved and required new framings.
Beyond these cognitive benefits for users, narrative-first design provided unique system-level benefits by making reasoning evolution visible and trackable. Externalizing thoughts in narrative form allowed participants to monitor how their assumptions shifted and revisit earlier claims. This highlights how narrative construction serves dual purposes: supporting human reasoning while providing systems with insight-level data that traditional interaction logs cannot capture (Ragan et al., 2015; North et al., 2011).
@@ 8.2. Redirecting Cognitive Effort and Supporting Productive Friction
Most interactive systems aim to reduce user effort under the assumption that easier interactions lead to better outcomes. However, cognitive load theory distinguishes between productive cognitive effort that supports learning and extraneous overhead that impedes it (Kalyuga, 2011; Sweller, 1994). Rather than simply minimizing effort, our findings suggest the importance of redirecting cognitive resources toward valuable reasoning activities. Our findings highlight that participants using NS reported significantly lower mental demand, effort, and frustration while engaging in more complex reasoning. For example, they generated 2.3 times more insights and conducting 2.7 times more reflection events. Also, they engaged in reflection throughout the whole process instead of only towards the end. This pattern suggests that NS reshaped how participants allocated cognitive resources during analysis.
@@ 8.3. Design Implications for Mixed-Initiative Analytical Systems
Effective divergence depends on helping users engage with multiple alternatives, rather than merely producing superficial options (Suh et al., 2024; Pu et al., 2025). Our bidirectional narrative–evidence coupling illustrates this. Features like Show View and Capture let users test a narrative hunch with visualizations, then save what they discover back into their reasoning. This cycle helps alternatives develop enough depth to support later reflection. While not every system needs a narrative-first interface, the principle applies broadly: mixed-initiative systems can support divergence by enabling exploration of multiple paths that help develop competing explanations from initial hunches into well-reasoned, evidence-backed alternatives.
@@ 8.4. Supporting Reasoning Evolution Across Domains
The principles underlying NS address a challenge that extends beyond data analysis: how to support reasoning that develops over time. A clear example is intelligence analysis, where analysts must synthesize fragmented information into coherent assessments while remaining open to contradictory evidence. NS suggests potential to track reasoning from initial hypotheses to evidence integration and final judgments, helping analysts maintain clear records of their reasoning and remain open to alternative explanations on ambiguous problems (Pherson and Heuer Jr, 2019; Heuer, 1999). Similar needs arise in policy development, where decision-makers must manage unresolved questions, weigh competing factors, and show how recommendations evolve from evidence (Hansen and Rieper, 2010). In research contexts, tools could likewise benefit from insight timelines that track the progression of hypotheses across literature reviews, experiments, and analysis.
@@ 8.5. Ethical Implications for Analytical Integrity
AI-assisted analysis systems raise concerns about analytical integrity when algorithmic decisions become invisible to users (Ha et al., 2024). We discuss three areas where NS’s design could unintentionally affect reasoning quality: algorithmic bias in view generation, narrative framing effects, and system steering through workflow design.
@@
While NS reduces cognitive overhead, the workflow design creates structural pressures. Prior work identifies steering challenges in AI-assisted analysis and proposes task decomposition to address them (Kazemitabaar et al., 2024). Narrative-driven exploration faces similar challenges at the workflow level and introduces friction through explicit actions for generating views, capturing insights, and creating branches. However, this friction occurs at execution points rather than validation points. Users must independently recognize when alternatives are needed or when reasoning should be questioned, and the system’s editorial choices during synthesis remain implicit. While NS makes branching and reflection visible workflow components, these depend on users initiating them. Future systems could explore friction-induced interventions that prompt alternative exploration when analytical patterns suggest narrowing, or make the system’s synthesis criteria more transparent.
@@ 9. Limitations and Future Work
This design choice means our findings may not generalize to high-stakes analytical contexts where correctness and bias have serious consequences (Wall et al., 2017; Li et al., 2025). In domains such as medical diagnosis (Saposnik et al., 2016) or policy evaluation (Korteling et al., 2023), reducing confirmation bias and ensuring analytical rigour are critical. While our study suggests NS supports broader exploration and deeper reflection, we cannot claim these process improvements translate into reduced bias or increased accuracy without targeted evaluation in such contexts.
@@
NS creates scaffolds for divergence through branching and alternative exploration but relies heavily on user-driven reflection to synthesize and converge insights. Without active reflection, participants may end up with scattered and unproductive exploration. Our analysis supports convergence through reflection was necessary for productive outcomes. However, systems could better support convergence processes through automated synthesis suggestions or propagation mechanisms that maintain coherence across alternative reasoning paths as users refine and integrate scattered explorations (Suh et al., 2025). This could include developing mechanisms to detect when users have sufficient alternatives for meaningful comparison and suggesting synthesis strategies.
@@ A.1. Semantic-Aligned View Generation Pipeline
Narrative propositions and visualization descriptions are then normalized into shared tags (i.e geo=, topic=, time=, and intent=) and embedded into a joint semantic space using a bi-encoder architecture (Wang et al., 2023). The joint space consolidates all components needed for alignment. It contains the narrative proposition templates, LLM-generated visualization captions, normalized tags, the expected data formats from each visualization specification, and pre-stored SQL queries produced by a text-to-SQL model. By combining narrative and visualization in one embedding space, the system maintains semantic consistency across both directions of investigation.
@@ A.4. Prompt Used for Interaction Capture
* If multiple views are relevant, you may combine two, but keep the sentence short.
* Ignore hovers that do not reveal new numbers.
@@ A.6. Prompt Used for Inquiry Labels
* text: short paraphrase or quote of the answer
* confidence: "high" (asserted/verified), "medium" (tentative), "low" (weak hint)
@@
* text: brief summary (data reference, mechanism, pattern, or comparison)
* basis: one of "data" | "mechanism" | "pattern" | "comparison" | "other"
@@
* suggested_by: Target issue was prompted as a follow-up/side question by this issue.
* generalized_from: Current issue is a broader abstraction of a more specific earlier issue.
* specialized_from: Current issue is a narrower/drilled-down version of a broader earlier issue.
* replaces: Current issue supersedes or rephrases an earlier issue (clearer scope/time/geo).

### v2

@@ 8.1. The Narrative-First Paradigm: Supporting Natural Reasoning Through Writing
Our findings reveal how narrative construction can function as a primary reasoning interface during data analysis, aligning with narrative psychology’s view that storytelling structures how people make sense of complex information (murray2003narrative; bruner1991narrative). Rather than treating writing as post-analysis documentation (zhao2021chartstory; lee2015more), participants used NS not merely to document completed insights, but as a thinking medium, a space where vague hunches evolved into precise, evidence-backed claims through the act of writing.
A key aspect of this narrative-as-thinking-medium approach was adapted to how participants actually think. This paradigm accommodated diverse analytical entry points and supported continuous interpretive refinement throughout exploration. Unlike visualization-first tools that require users to translate thoughts into structured queries (setlur2016eviza; gao2015datatone), NS accommodated diverse entry points—whether specific hypotheses or exploratory questions. This addresses the ”gulf of envisioning” (subramonyam2024bridging): the difficulty of translating vague intentions into system-actionable inputs. Participants could write underspecified hunches, generate visualizations, then iteratively refine their narrative to pursue new directions—each refinement triggering updated views without breaking flow. This fluidity reduced friction not just during early-stage exploration (wongsuphasawat2019goals; kandel2012enterprise), but throughout the entire sensemaking process, as participants’ understanding evolved and required new framings.
Beyond these cognitive benefits for users, narrative-first design provided unique system-level benefits by making reasoning evolution visible and trackable. Externalizing thoughts in narrative form allowed participants to monitor how their assumptions shifted and revisit earlier claims. This highlights how narrative construction serves dual purposes: supporting human reasoning while providing systems with insight-level data that traditional interaction logs cannot capture (ragan2015characterizing; north2011analytic).
@@ 8.2. Redirecting Cognitive Effort and Supporting Productive Friction
Most interactive systems aim to reduce user effort under the assumption that easier interactions lead to better outcomes. However, cognitive load theory distinguishes between productive cognitive effort that supports learning and extraneous overhead that impedes it (kalyuga2011cognitive; sweller1994cognitive). Rather than simply minimizing effort, our findings suggest the importance of redirecting cognitive resources toward valuable reasoning activities. Our findings highlight that participants using NS reported significantly lower mental demand, effort, and frustration while engaging in more complex reasoning. For example, they generated 2.3 times more insights and conducting 2.7 times more reflection events. Also, they engaged in reflection throughout the whole process instead of only towards the end. This pattern suggests that NS reshaped how participants allocated cognitive resources during analysis.
@@ 8.3. Design Implications for Mixed-Initiative Analytical Systems
Effective divergence depends on helping users engage with multiple alternatives, rather than merely producing superficial options (suh2024luminate; pu2025ideasynth). Our bidirectional narrative–evidence coupling illustrates this. Features like Show View and Capture let users test a narrative hunch with visualizations, then save what they discover back into their reasoning. This cycle helps alternatives develop enough depth to support later reflection. While not every system needs a narrative-first interface, the principle applies broadly: mixed-initiative systems can support divergence by enabling exploration of multiple paths that help develop competing explanations from initial hunches into well-reasoned, evidence-backed alternatives.
@@ 8.4. Supporting Reasoning Evolution Across Domains
The principles underlying NS address a challenge that extends beyond data analysis: how to support reasoning that develops over time. A clear example is intelligence analysis, where analysts must synthesize fragmented information into coherent assessments while remaining open to contradictory evidence. NS suggests potential to track reasoning from initial hypotheses to evidence integration and final judgments, helping analysts maintain clear records of their reasoning and remain open to alternative explanations on ambiguous problems (pherson2019structured; heuer1999psychology). Similar needs arise in policy development, where decision-makers must manage unresolved questions, weigh competing factors, and show how recommendations evolve from evidence (hansen2010politics). In research contexts, tools could likewise benefit from insight timelines that track the progression of hypotheses across literature reviews, experiments, and analysis.
@@ 8.5. Ethical Implications for Analytical Integrity
AI-assisted analysis systems raise concerns about analytical integrity when algorithmic decisions become invisible to users (ha2024guided). We discuss three areas where NS’s design could unintentionally affect reasoning quality: algorithmic bias in view generation, narrative framing effects, and system steering through workflow design.
@@
While NS reduces cognitive overhead, the workflow design creates structural pressures. Prior work identifies steering challenges in AI-assisted analysis and proposes task decomposition to address them (kazemitabaar2024improving). Narrative-driven exploration faces similar challenges at the workflow level and introduces friction through explicit actions for generating views, capturing insights, and creating branches. However, this friction occurs at execution points rather than validation points. Users must independently recognize when alternatives are needed or when reasoning should be questioned, and the system’s editorial choices during synthesis remain implicit. While NS makes branching and reflection visible workflow components, these depend on users initiating them. Future systems could explore friction-induced interventions that prompt alternative exploration when analytical patterns suggest narrowing, or make the system’s synthesis criteria more transparent.
@@ 9. Limitations and Future Work
This design choice means our findings may not generalize to high-stakes analytical contexts where correctness and bias have serious consequences (wall2017warning; confirmation2025). In domains such as medical diagnosis (saposnik2016cognitive) or policy evaluation (korteling2023cognitive), reducing confirmation bias and ensuring analytical rigor are critical. While our study suggests NS supports broader exploration and deeper reflection, we cannot claim these process improvements translate into reduced bias or increased accuracy without targeted evaluation in such contexts.
@@
NS creates scaffolds for divergence through branching and alternative exploration but relies heavily on user-driven reflection to synthesize and converge insights. Without active reflection, participants may end up with scattered and unproductive exploration. Our analysis supports convergence through reflection was necessary for productive outcomes. However, systems could better support convergence processes through automated synthesis suggestions or propagation mechanisms that maintain coherence across alternative reasoning paths as users refine and integrate scattered explorations (suh2025storyensemble). This could include developing mechanisms to detect when users have sufficient alternatives for meaningful comparison and suggesting synthesis strategies.
@@ A.1. Semantic-Aligned View Generation Pipeline
Narrative propositions and visualization descriptions are then normalized into shared tags (i.e geo=, topic=, time=, and intent=) and embedded into a joint semantic space using a bi-encoder architecture (wang2023improving). The joint space consolidates all components needed for alignment. It contains the narrative proposition templates, LLM-generated visualization captions, normalized tags, the expected data formats from each visualization specification, and pre-stored SQL queries produced by a text-to-SQL model. By combining narrative and visualization in one embedding space, the system maintains semantic consistency across both directions of investigation.
@@ A.4. Prompt Used for Interaction Capture
[…]
@@ A.6. Prompt Used for Inquiry Labels
[…]
@@
[…]
@@
[…]